OpenStudy (anonymous):

Find a polar equation of the hyperbola (x/11)^2−(y/10)^2=1. r^2=?

OpenStudy (dumbcow):

x = rcos y = rsin \[r^2 = \frac{12100}{100\cos^2 \theta - 121\sin^2 \theta}\]